OpenStudy (anonymous):

Verify the identity (1 + tan2u)(1 - sin2u) = 1

OpenStudy (owlcoffee):

you want to prove it?

OpenStudy (jdoe0001):

notice the trig pythagorean identities -> http://web.mit.edu/wwmath/trig/eq1.gif

terenzreignz (terenzreignz):

Just need to play around with this known identity:\[\Large \cos^2(x)+\sin^2(x) = 1\]

OpenStudy (anonymous):

Yeah, I don't understand identities very well, so how is it proven?

OpenStudy (jdoe0001):

can you solve \(\bf \cos^2(x)+\sin^2(x) = 1\) for \(\bf cos^2(x)?\) what do you get?

OpenStudy (anonymous):

But how do you get from cos and sin to tan? Like my problem up top?

OpenStudy (anonymous):

\[\cos ^{2}x = 1 - \sin ^{2}x\] but how does it get to \[\tan ^{2}\]?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Can you please explain how it works? @jdoe0001

terenzreignz (terenzreignz):

You seem panicked :) see what happens when the entire equation \[\Large \cos^2(x)+\sin^2(x) = 1\] is divided by \(\large \cos^2(x)\)... \[\Large \frac{\cos^2(x)}{\cos^2(x) }+ \frac{\sin^2(x)}{\cos^2(x)}= \frac1{\cos^2(x)}\]

terenzreignz (terenzreignz):

Obviously, this red part is just 1. \[\Large \color{red}{\frac{\cos^2(x)}{\cos^2(x)} }+ \color{blue}{\frac{\sin^2(x)}{\cos^2(x)}}=\color{green}{ \frac1{\cos^2(x)}}\] This blue part is \(\large \tan^2(x)\)\[\Large \color{red}{1 }+ \color{blue}{\frac{\sin^2(x)}{\cos^2(x)}}=\color{green}{ \frac1{\cos^2(x)}}\] And this green part is just \(\large \sec^2(x)\) \[\Large \color{red}1+\color{blue}{\tan^2(x)}= \color{green}{\sec^2(x)}\]

OpenStudy (anonymous):

thank you!

terenzreignz (terenzreignz):

No problem.